A Complete Bibliography of Publications in Biometrika for the decade 1930–1939

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.04

Title word cross-reference

(χ^{2}, P) [Pea32b]. (n < 100) [Pea32a]. − log_{10} λ_{n} [Dav34a]. 4 [McK33a]. β_{2} [Gea35a, McK33a, Pea35a]. χ [JW39]. χ^{2} [Coc38, El 36, Hal37, Hal38b, Hal39a, Hal39b, Les33, NP31]. k [PW33, Wel35]. K_{τ_{1},τ_{2}}(X) [Pea33d, Rig33]. λ_{H_{1}} [WT37]. n [Hal38b, Hal39b, MP33, McK35, Pea33c]. P [Les33]. P_{λ_{n}} [Dav34a]. ϕ^{2} [Pea33e]. ψ^{2} [Dav39]. \sqrt{beta_{1}} [McK33b]. √β_{1} [Wil35]. T_{m}(x) [PSD32]. T_{τ_{1},τ_{2}}(x) [Rig33]. w′ [Gea35a]. w_{n} [Pea35a]. x [Dav33]. z [Cha38, Law38, Law39, Pea31d]. Z_{1} [Pea31c]. Z_{2} [Pea31c].

-distribution [JW39]. -Fold [Hal38b, Hal39b].

A. [Ano32e, Chu32]. Abortion [Pea30a]. Absence [San38]. Acaulis [Hal38c]. Account [Til33]. Accuracy [El 36, Mar32]. Actual [Nai36a]. Actuarial [Ste31]. Adjustments [Eld33]. Advantages [Yat39]. Affinities [You31]. Again [Cam34]. Age [And33, Les33]. Agricultural [Pea38c, Yat39]. Aiken [B.36]. Albanians [Til33]. alle [Har34]. Ambulatory [SM32]. American [Sus33]. American-Born [Sus33]. Among [Wel35]. Amphibia [Ter31]. Analysis [Hal39c, Pea31a, PW33, Pit38, Wel35, Wil32, Wis38]. Anglo [LY35, Mün36]. Anthropology [Mor39b]. Anthropometric [Mah30]. any [Pit38, Pit39a]. Applicability [Cha38]. Application [Hen37, Mul31, Nai39, PW36, Yul39]. Applications [PSD32, Pea33d]. Applied [GH37, Pit38]. Appreciation [Pea36a, Pea38a]. Appropriate [PW33]. Approximate [Hal38a]. Approximation [El 36]. Archaeological [Ris39]. Arise [PS31]. Arising [Tan38]. Arms [Mün36]. Arrangements [Jef39, NP38, Stu38, Yat39]. Arterio [GR38]. Arterio-Sclerosis [GR38]. Arthur [B.36]. Article [Bat34]. Ascertaining [Nai36a]. Asia [Von31]. Asiatic [WM32]. Asien [Von31]. Aspects [Pea36a, Pea37b, Pea38a, Pea38c, Tho32]. Asymmetry [Woo31]. Asympotic [Dav34b]. Asymptotic [Dav33]. Aurel [Mor38]. Australia [Wun39]. Authorship [Yul39]. Avon [BY35].

B [Coc38, Mor35]. Back [Ano30a, Ano31b, Ano31a, Ano32a, Ano32b, Ano33a, Ano33b, Ano34a, Ano34b, Ano34c, Ano35a, Ano35b, Ano36a, Ano36b, Ano37a, Ano38b, Ano38a, Ano39b, Ano39a]. Bacon [Wis38]. Bacterial [Gor39]. Badari [Mor35]. Balanced [NP38, Stu38]. Balancing [Pea38c]. Based [HM36, Mor35, Mor37, WM32, Yul38b]. be [Mar34, Pit38]. Bedfordshire [DY33]. bedingten [Mah35]. Been [Pea33c]. Being [Pea35b]. Beitrag [Mah35, Von31]. Bessel [McK32, PSD32, Pea33d, Rig33]. Beta [Mul31]. Beta-Function [Mul31]. Betas [Dav32]. Between [Gea35a, Hoj33, Hot36, McK35, Mor39c, NP38, PP32, Ste34, Stu38, Wel38b, Pea31c]. Bewegungen [Mah35]. Bi [Ano30d, Fin38, Gar33]. Bi-Variate [Ano30d, Fin38, Gar33]. Bidford [BY35]. Bidford-on-Avon [BY35]. Binomial [Ayy34b, CP34, Ney38]. Biological [Yat39]. Biometric [Cle37, MCA36, WM34, Woo37, Ter31]. Biometrician [B.36]. Biometrika [Pea35e, Ano30c, Ano32e, Ano33e]. Biometrische [Ter31]. Births [Bel33]. Bitches [PP31]. Bivariate [Bra33, Cha32, Fie32, Hir37, Mor39c, Pea31c, Pre30, Yul38b]. Blocks [Wel37]. Boas [MS36]. Bodily [BD35]. Body [Sus33]. Bone [Woo37]. Bones [Mün36, PW35a]. Book [B.36, Dav38, H.39, Lun38, S.39]. Born [Sus33]. Botanist [B.36]. Breeding [Tan38]. Bright [GR38]. Britain [vB36]. Broad [Pea35b]. Build [Sus33]. Burr [B.36]. Burwell [LY35]. Busts [PM34].

C [B.36, H.39, SBL79]. calculation [JW39]. Carota [Bat33, Bat34]. Cartesian [Pea33b]. Case [Cha32, CP34, PS31, PP32]. Cases [HM36, Pea31a, Yul39]. Categories [Pea35b]. Causes [Kar33]. Cell [Cha32]. Certain [Gea33, Han34, Hey38, Kar33, LY35, Mah30, Mul31, PP32, Ste34, Tan38, Wil32, Wis30, Wiś35, Wis36]. Changes [BD35, Ush32]. Character [PP32]. Characteristic [Yul39]. Characteristics [Mar34, Ter31].

Characters [Les33, Mor39b, PP31, PW35a]. Children [BD35, Sus33]. Chrysanthemum [Bat35]. Church [SM32, Chu32]. Cicuta [Bat36]. Class [Hal38a, Kol35]. Classification [Mor39b, WM32]. Cleft [Coc36]. Clover [PW35b]. Coefficient [Dav37, Gar33, KKS39, Pep32, Rid32, WKE31]. Coefficients [Bra33, Mar34]. Collections [Wun39]. Combining [Pea38b]. Comments [SBL79]. Commonwealth [Wun39]. Comparative [MC38, Yat39]. Comparison [LY35, Mar32, NP38, Pea35a, Stu38, Wis33]. Comprehensive [Bis39]. computer [Tei65]. Concerning [Pit39c]. conditional [Mah35]. Confidence [CP34]. Configurations [Tho32, THM^{+}32]. Connected [Rig33]. Constancy [Bat35]. Constructing [Kar33]. Contingency [Ano34g]. Continued [Mul31]. Continuous [Pit39a]. Contribution [Cle37, Die34, GR38, Mor37, Pea39a, Mah35, Die34, Von31]. Contributions [St.32]. Converse [Cam32]. Coordinatograph [Pea33b]. Correction [Hal39a, Hal39b, Law39]. Corrections [Mar34, Pea33a, Pea35b]. Correlation [Bra33, Dav37, Die34, Gar33, Gea35a, Ken38, KKS39, Pea31d, Pit39b, Rid32, Rie32, Wic33, WKE31, Wis32, Lun38]. Corrigenda [Ano48a, Ano48b]. Covariance [Hir37]. Covariances [Bis39]. Crania [MH31, Mor39a, SM32, vB36]. Cranial [EW32, Pea33b, Ris39, VM38, WM32, Wun39]. Craniologist [Pea33b]. Craniology [Mor37, vB36, Von31]. Criteria [Pea35a]. Criterion [Dav39, PS36, WT37]. Cromwell [PM34]. cumulants [JW39]. Curves [Ano33g, Eld33, Lun38, PS31, PP32, Pea33a].

D [Mor35, Pea39b]. Dacryocystitis [Coc36]. Data [San38]. Daucus [Bat33, Bat34]. Dealing [MS36]. Death [Kar33]. Deducation [Ney38]. Degrees [Hal38b, Hal39b]. Dependence [Ste34]. Derivation [Wis30]. Derry [Mor35]. Description [Mor38]. Design [Yat39]. Determinations [Wis38]. Determined [Mar34]. Determining [Ano30d, Pea33c, Pea34b]. Development [Pea30b]. Developmental [BD35]. Deviation [Ano48a, Ano48b, Gea35b, Gea36, New39, Pea31d, SBL79]. Deviations [Kon30]. Diameters [Wag35]. Difference [Bra33, McK35, Mor39c, Nai36b, Wel38b]. Differences [EM31]. Dilution [Gor39]. dimensional [Mah35]. Direct [Pea39b]. Discriminant [Wel39a]. Discussion [Pea32b]. Disease [CMR39, GR38]. Disputed [Yul39]. Distribution [Cha38, Coc38, Fie32, Fin38, Gar36, Hal37, Hal39a, Hir37, KKS39, Kon30, Le 31, Led39, McK32, McK33b, McK33a, MP33, McK35, Nai39, New39, NP31, Pea30a, Pea32a, Pep32, Per33, Rid32, Wel39b, WT37, Wil35, Wis32, JW39, Tei65, Dav39]. Distributions [Dav32, EW32, Hal38a, Han34, Mar34, Mor39b, Pea32c, Wis33, Yul38b, Mah35]. Dodder [PW35b]. Double [Pea33d, Rig33]. Dr [MS36, Pea35c, Pea35d]. Dr. [Mor39a]. Drawn [Hir37, Pea33c, Per33]. Dunn [H.39]. Dunstable [DY33]. Duration [Fie31a, PP31]. During [BD35, Bat35]. Duweir [Ris39]. Dynasty [Col33, Woo30].

East [Von31]. Editorial [Pea45]. Effect [Pea38c]. Efficiency [PS36]. Egyptian [Mar36, Mor35, Woo30]. Elderton [Lun38, Pea33a]. Element [Dav33, Dav34b]. Eliminated [Kar33]. Empirical [And33]. Endocranial [Wag35]. Entropion [Ush32]. Environment [CMR39]. Epicanthus [Pea35c, Ush35]. Equal [KKS39]. era [Tei65]. Ereignisse [Lud34]. Error [Nai36b]. Eruption [HM36]. Eskimo [Mor37]. Estimate [New39]. Estimates [Fin38, Hir37, Wel39b, Wis33]. Estimating [Bea39, Gor39]. Estimation [Pea37a, Pea39b, Pea39a, Pit39a]. étude [Die34]. Eulerian [PP35]. Evaluation [Mul31, PP35]. events [Lud34]. Evidence [Tho33]. Exact [Coc38, Hal37, Hal39a]. Examination [MS36]. Examined [Pre30]. Excavated [MH31, Mor38, Ris39]. Excavations [DY33]. Exceeds [Tho33]. Expectations [Hal37, Hal38b, Hal39a, Hal39b]. Expedition [Ris39]. Experiment [Pea38c, Pep32]. Experimental [Hey38, Pea32b]. Experiments [Yat39]. Expressed [New39]. Expression [Gea33]. Extreme [McK35]. Eye [TE35]. Eyelids [Ush32].

Facial [WM34]. Factor [Cam32, Cam34]. Family [Coc35, Nus38]. Feet [Coc36]. Fiducial [CP34, Gar36]. Field [Bea39, NP38, Stu38]. First [BD35, Hal38b, Hal39b]. Fishberg [MS36]. Fisher [Law39, Law38]. Fit [Dav39, GH37, Hal37, Hal39a, Pea32b, Pea34b, Pea38b]. Fitting [Ano33g]. Five [S.39]. Flatness [WM34]. Flowering [Bat35]. Flowers [Bat35]. Fold [Hal38b, Hal39b]. Form [Ber38, Mor39b, Pit39a]. Formula [Bay38, Wiś35, Wis36]. Formulae [Ayy34b, Dav33, Dav34b, Mar32]. Four [McK33b]. Fourth [Dav33, Dav34b]. Fractions [Mul31]. Franz [MS36]. Frau [PW36]. Freedom [Hal38b, Hal39b]. French [Die34]. Frequencies [Cha32, Yul38b]. Frequency [Ano33g, Cha32, EW32, Hal38a, Han34, Mar34, Pea30a, PS31, Pea32c, Pre30, Lun38]. Front [Ano30b, Ano31d, Ano31c, Ano32c, Ano32d, Ano33c, Ano33d, Ano34d, Ano34e, Ano34f, Ano35c, Ano35d, Ano36c, Ano36d, Ano37b, Ano38d, Ano38c, Ano39d, Ano39c]. Function [McK32, Mul31, Nai39, PSD32, Pea33d]. Functions [Gea33, Rig33, Wel39a, Wic33]. Further [Dav34a, Hoj33, Kar33, NP31, Pea30b, PSD32, PW35a, St.32].

Gauss’ [Bay38]. Geary [Pea35a, SBL79]. Geiringer [PW36]. General [Cam34, Gea33, Mor39b, Tho32]. Generalization [Law38, Law39]. Generalizations [Wil32]. Genetics [H.39]. Gentile [Pea36c]. Geometric [Tho32, THM^{+}32]. Geometry [Pea33b]. George [B.36]. German [Har34, Lud34, Mah35, Ter31, Von31]. Gini [Nai36b]. Given [Dav34a, Nai36a, Pea33e, Pit39a]. Gives [Pea33e]. Goodness [Dav39, GH37, Hal37, Hal39a, Pea32b, Pea34b, Pea38b]. Gordon [Pea39b]. Gortner [B.36]. Growth [BD35, Wis38]. Growth-Rate [Wis38].

Hair [TE35]. Haldane [Coc38]. Hand [Coc36]. Handling [Pea34a]. Hare [Coc36]. Harris [B.36]. Having [Pea33c]. Head [PM34]. Health [S.39].

Hereditary [Ush32, Ush35]. Heredity [Moo35]. Heterostylism [Hal38c]. High [PP35, Wis30]. Historical [Ney38]. History [GR38, Tei65]. Hojo [Rom33]. Homogeneity [Bis39, Wel38a]. Human [Cle37, Moo35, MCA36, Mor38, PW35a, Ris39, Woo31, Woo37]. Hypergeometric [Dav33, Dav34b, Dav33]. Hypergeometrical [Ayy34a]. Hypotheses [Kol35, Pit39c]. Hypothesis [Dav39, WT37]. Hythe [SM32].

II [PP32, Pea38c, THM^{+}32]. III [Pit38, Wic33]. Illustrated [CP34, Hey38, San38]. Illustration [Ano30c, Ano32e, Ano33e, Ano36e, Dav34a, El 36, Les33, Pea38c]. Illustrations [Pea33b, Pre30]. Impaction [HM36]. Important [Kol35]. Incidence [MC38, THM^{+}32]. Including [LY35]. Incomplete [Ayy34a, Mul31, PP35]. Independent [New39, Pea32c, Pea33e, Pea38b]. Index [Fie32]. Indian [VM38]. Indices [Wag35]. Individual [PP32, PW35a]. Individuals [PP32]. Infinite [Wis33]. Inflorescences [Bat36]. Influence [Bat36]. Information [Ano31e, Ano32g, Ano33h, Ano34h, Ano35e, Ano36g, Ano38e, Ano39e]. Inheritance [RDM39]. Inquiry [Pea38c]. Insects [Bea39]. Instrument [Wag35]. Integral [Gar33, Pea33c, Pea38b]. Integrals [Mul31, PS31, PP35, Rom33]. Internal [Wag35]. Intervals [PP32, San38]. Intraracial [Mor39b]. Introduction [Dav38]. Introductory [Til33]. Invariant [Wis33]. Invariants [Wis33]. Inverse [Pea39b]. Investigation [Mor39b, PW35a]. Investigations [MS36]. Iran [Mor38]. Irregular [San38]. Islands [THM^{+}32].

J [Dav38, Mor39a, Eld33, Pea33a]. J-Curves [Pea33a]. J-Shaped [Eld33]. J. [Coc38, B.36]. Japanese [Sus33]. Jewish [Pea36c]. Jews [MS36].

K.C.I.E. [Mor38]. Karl [Ano36f, Nai36a, Ney38, Pea36a, Yul38a]. Kerma [Col33]. Kn [Ano33e]. Known [Pea33c]. Kolodziejcyk [Pea35e]. Kraniologie [Von31].

L [H.39]. L. [Bat33, Bat34]. Lachish [Ris39]. Langerhans [THM^{+}32]. Latin [Wel37]. Laws [Nai39]. Lectures [Yul38a]. Legs [Mün36]. Length [Yul39]. Lengths [Mün36]. Leonard [SM32]. Letty [Ano33e]. Leucanthemium [Bat35]. Life [Kar33, Pea36a, Pea38a]. Light [Pre30]. Ligulate [Bat35]. like [Mah35]. Likelihood [Hen37, Pea36b, Pea37a, Tho33, Wel39b]. Limits [CP34, Dav37, Gar36, Nai36a, Pea32a]. Lip [Coc36]. Litter [PP31]. Location [Pit39a, Pit39c]. Logischen [Ter31]. London [You38]. Long [Mün36]. Low [PS31].

Maculata [Bat36]. Mainly [Moo35]. Malar [Woo37]. Man [HM36, Les33, McM39, Mün36, WM34]. Mandible [Cle37, MCA36]. Mandibles [Mar36]. Mandibular [HM36]. Marked [Nai36a]. Markoff [Ano32e]. Marston [Ris39]. Masks [PM34]. mass [Mah35]. Massenpunkten [Mah35]. Material [Mor39b]. Mathematical [Dav38, Mar36]. Matter [Ano30a, Ano30b, Ano31b, Ano31a, Ano31d, Ano31c, Ano32a, Ano32b, Ano32c, Ano32d, Ano33a, Ano33b, Ano33c, Ano33d, Ano34a, Ano34b, Ano34c, Ano34d, Ano34e, Ano34f, Ano35a, Ano35b, Ano35c, Ano35d, Ano36a, Ano36b, Ano36c, Ano36d, Ano37a, Ano37b, Ano38b, Ano38a, Ano38d, Ano38c, Ano39b, Ano39a, Ano39d, Ano39c]. Maurice [MS36]. Maximum [Hen37, Pea36b, Pea37a, Wel39b]. May [Pit38]. McKay’s [Ano33f]. Mean [Ano48a, Ano48b, Gea35b, Gea36, McK35, Nai36b, Pea31d, PP32, Rom33, SBL79, WKE31]. Means [Wel38b]. Measure [Ano34g, Ken38]. Measurements [BD35, EW32, HM36, Mah30, Mor35, MS36, Mor37, Til33, VM38, WM32]. Measures [Ste34]. Measuring [Wag35]. Median [Hoj33]. Medical [GR38]. Members [Nai36a]. Memoir [Chu32, Pea35b, Pea35d]. Mendelian [GH37, Hal39c, RDM39]. Merkmale [Ter31]. Messrs [Pea35e]. Method [Cha32, Gor39, Hen37, Hey38, Nai36a, Pea33b, Pea33c, Pea34b, Pea36b, Rom36]. Methods [Bea39, Kar33, Les33, Mar36, Mor39b, PW33, Pea37a, Pea39b]. Metrical [Mor39b]. Mice [GH37]. Microscopic [Tho32, THM^{+}32]. Minimum [Pea33e]. Miss [Mor35]. Molar [HM36]. Moment [Mar34, Nai39, Pep32, WKE31, Wis33]. Moments [Ano48a, Ayy34a, Ayy34b, Coc38, Eld33, Gea33, Gea36, Hal37, Hal38b, Hal39a, Hal39b, Ney38, Pea33a, Pea36b, Rom36, Wis30]. Monkey [THM^{+}32]. Morpho [Ter31]. Morpho-Logischen [Ter31]. morphological [Ter31]. Morphometric [PW35a]. Mortality [MC38]. movements [Mah35]. Mr [Ano33f, Pea33a, Pea35b, Pea35a]. Multiple [WKE31]. Multivariate [Bis39].

N [Mor35]. Naga [KM33]. Natural [Hal38c]. Nature [Mor39b, Pea31c]. Necessarily [Dav34b]. Nephritis [GR38]. Neyman [Dav39]. Nine [Mor38]. Ninth [Woo30]. Non [Han34, Hey38, Pea31a, Rie32]. Non-Normal [Han34, Hey38, Pea31a, Rie32]. Normal [Ano30d, Ano48a, Bra33, Fie32, Fin38, Gar33, Gea33, Gea36, Han34, Hey38, Hir37, Hoj33, Led39, McK33b, McK33a, McK35, Mor39c, New39, Pea31a, Pea31c, Pea32a, Pit39b, Rie32, Rom33, Wil35, Wis30, WKE31, Yul38b]. Normality [EW32, Gea35b, Pea30b, Pea31b]. Normalization [Hal38a]. North [Til33]. Note [Ano33g, Ano33f, Ayy34a, Ayy34b, Ber38, Chu32, Coc38, Dav37, Gea35a, Hal39c, Hoj33, Les33, MP33, Mor39a, Nai36a, Ney38, NP38, Pea31b, Pea33a, Pea35b, Pea35c, Pea35d, Pea39b, Pea39a, Pea45, Pit39b, PW36, Rom36, Wel39a, Wil35, Wis32]. Notes [NP31, Yul38a]. Nubia [Col33]. Number [Bat35, KKS39, Nai36a]. numbers [Nai38]. Numerical [PP35, Pre30]. Nutrition [Wis38].

O [B.36]. Observation [McK35]. Observations [PS36]. Observed [RDM39]. Obtained [HM36]. Occur [KKS39]. Oceania [vB36]. Oliver [PM34]. One [BD35, Mar34, Tho33, Woo30]. Order [PP35, Wis30]. Ordinates [Bay38, Cha32]. Ost [Von31]. Ost-Asien [Von31]. Other [LY35, PP31, Ris39, Wun39]. Otto [B.36]. Outlying [PS36].

Painted [PM34]. Palate [Coc36]. Palestine [Ris39]. Palin [Lun38, Pea33a]. Pall [Ter31]. Pancreas [THM^{+}32]. Paper [Ano33f, Coc38, Nai36a, NP38, Pea35c, SBL79]. Papers [Pea35e]. Parameters [Pit39a, Pit39c]. Parent [Pea32c, Pea33c, Pea33e]. Part [Mor39b, PP32, Pea33d, Pea36a]. Partial [Cha32]. Pearson [Ano36f, Nai36a, Ney38, Pea36a, Wiś35, Wis36, Yul38a]. Pedigree [Bel33]. Pedigrees [Ush35]. Percentage [Pea32a]. Photographs [Til33]. Physical [Les33]. Piebald [Coc35, Nus38]. Pig [Wis38]. Pigmentation [TE35]. Pinnatifidum [Bat35]. Pitman [Pea39a]. Planes [Pea33b]. Plant [Tan38]. Play [Fie31a]. Plots [NP38, Stu38]. Plural [Bel33]. Point [Ayy34b]. Points [NP38, Mah35]. Poisson [Cha38, Gar36]. Pollaczek [PW36]. Pollaczek-Geiringer [PW36]. Population [Bea39, Fie32, Fin38, Gar33, Hoj33, Led39, Mor39c, Nai36a, New39, Pea31c, Pea32a, Pea32c, Pea33c, Pea33e, Pit39a, Rom33, Wel38b, Wil35, Wis30, WKE31, Wis33]. Populations [Gor39, Hal38c, Hey38, Hir37, Pit38]. Portrait [Ano30c, Ano32e, Ano33e]. Portraits [PM34]. Position [Bat36]. Powers [EM31, PS31]. pp [Pea35c]. practice [Har34]. Praxis [Har34]. Preceding [Hal39c]. Predicting [HM36]. Predynastic [Mor35]. Pregnancy [PP31]. Preliminary [HM36, WM32]. Previously [Mor37, VM38]. Primrose [Hal38c]. Primula [Hal38c]. Principles [PW35b, H.39]. prior [Tei65]. Probability [Ano48b, Fie31b, Gar33, PS31, Pea32c, Pea33c, Pea38b, Tho33, Dav38]. Probably [Pea33c]. Problem [Ano32f, Fie31b, HM36, Pea37b, Pea38c, Pea39b, St.32]. Problems [Bis39, Mor39b, Pea33d, Tan38, Wel35]. Proceeding [Cha32]. Product [Wis30]. Prof. [Rom33]. Professor [Ano32f, Ano34g, Mor35, MS36, Pea39a]. Progress [S.39]. Projectometry [Tho32]. Properties [Pea31d, Tho32, THM^{+}32, Yul38b]. Property [Rom33]. Proportions [BD35]. Prose [Yul39]. Published [Mor37, VM38]. purposes [Har34].

Quadratic [Bay38]. Quadrature [Mar32]. Quadrilateral [Dav32]. Quartiles [Hoj33].

R [Chu32]. R. [Pea39b, SBL79]. Races [VM38, WM32]. Radiographs [HM36]. Rana [Ter31, Ter31]. Random [Jef39, NP38, Pea33c, Stu38, Mah35, Nai38]. random-like [Mah35]. Randomization [Pea37b, Pea38c]. Randomized [Wel37, Yat39]. Randomness [Dav34a]. Range [MP33, New39, Pea32a]. Ranges [Rom33].

Rank [Ken38, KKS39]. Ranked [PP32]. Rankings [KKS39]. rare [Lud34]. Rate [Wis38]. Ratio [Ano48a, Fin38, Gea35b, Gea36, Hir37, Per33, Wis32]. Ratios [RDM39]. Really [Pea32c]. Recently [MH31]. Recorded [San38]. Records [GH37]. Recovery [San38]. Rectangular [Per33]. Recurrence [Ayy34b, Rig33]. Reference [HM36, Mar36, Mul31, Mün36]. Regression [Pea31d, Wel35]. Rejection [PS36]. Rejoinder [Wis36]. Relation [Hoj33, PP31, Rig33, TE35]. Relations [Hot36]. Relationship [Pea31c, PM34]. Relationships [Pea36c]. relevance [Tei65]. Remaining [Ano30d]. Remains [Ris39, Wun39]. Remarks [Ano34g, Dav34a]. Repeated [BD35]. Research [Ris39]. Researches [Ste31]. Review [B.36, Dav38, H.39, Lun38, S.39]. Ridibunda [Ter31]. Rietz [Ano32f]. Rosendahl [B.36]. Ross [B.36]. Roumanian [Per34]. Routine [PW35b]. Rural [MC38].

S [Coc38]. Saint [SM32]. Salientia [Ter31]. Same [Pea32c]. Sample [Fin38, McK35, Mor39c, Nai36a, Pea31d, Pea33c, Pea33e]. Samples [Ano48a, Bra33, Gar33, Gea33, Gea36, Hir37, Hoj33, Le 31, McK33b, McK33a, MP33, McK35, New39, Pea31c, PS31, Pea32a, Pea32c, PW33, Per33, Pit38, Rid32, Rom33, Tho33, Wel35, WKE31, Wis33]. Sampling [Hey38, Kon30, Led39, Pep32, St.32, WT37, Wil35, Wis30, Nai38, Tei65]. Saxon [LY35, Mün36]. Scale [And33, Pit39a, Pit39c]. Scedasticit [Pea31d]. School [San38]. Science [Ste31, Har34]. Sclerosis [GR38]. Scottish [You31]. Season [Bat35]. Second [Bat34, Coc35, WKE31]. Section [Pea35d]. Sedment [Woo30]. Seed [PW35b]. Segregation [GH37]. Segregations [Hal39c]. Selection [Led39]. Seltenen [Lud34]. Semi [Wis33]. Semi-Invariant [Wis33]. Semi-Invariants [Wis33]. Sentence [Yul39]. Sentence-Length [Yul39]. Series [Ano30c, Ano32e, Ano33e, Ayy34a, Dav33, Dav34b, LY35, Mar36]. Sets [Hot36]. Seventy [Woo30]. Seventy-One [Woo30]. Sex [MC38]. Sexing [Mar36]. Shaped [Eld33]. Shrinkage [Les33]. Significance [Bra33, Mor39c, Pea38b, Pit38, Wel38b]. Silhouette [Per34]. Simometers [Pea34a]. Simple [Rie32]. simulation [Tei65]. Sinnott [H.39]. Sir [Mor38]. Six [Hal38b, Hal39b]. Sixty [BD35]. Sixty-One [BD35]. Size [BD35, Nai36a, PP31, Pea33c]. Skeletal [Wun39]. Skeleton [WM34]. Skeletons [Mün36]. Skew [PP32, Pre30]. Skin [Ush32]. Skull [KM33, Pea33b, PW35a, Wag35, Woo31, You31, You38]. Skulls [BY35, Col33, DY33, LY35, Mor35, Mor38, Woo30]. Small [Hal37, Hal38b, Hal39a, Hal39b, Hoj33, Le 31, PS31, Rid32]. Smooth [Dav39]. Some [NP38, Pea31d, Pea33b, Pea36a, Pea37b, Pea38c, Rom33, Ste31, Wel35, Yul38b, Hal38b, Hal39b, Pea38a]. South [Til33]. Spearman [Cam32, Cam34, KKS39]. Special [HM36, Mar36, Mul31, Mün36]. Spitalfields [MH31]. Squares [Wel37, Yul38b]. Standard [Ano48a, Gea35b, Gea36, Kon30, Nai36b, New39, Pea31d, Pea33b]. Standardized [SBL79]. Start [Mar34]. States [VM38]. Statistical [Bat33, Bat34, Kol35, Mah30, Mor39b, Pea33d, PW33, PS36, PW35b, Ste34, Tan38, Yul39, Har34]. Statistician [Pea39c]. Statistics [GR38, Nai39, PSD32, Ste31, Yul38a, Lud34]. Statistik [Lud34]. Statistische [Har34]. Steffensen [Ano34g]. Stein [Mor38]. Stoessiger [Mor35]. Strength [TE35]. Structure [Bat36]. Student [McM39, Pea39c, Per33]. Student’s [NP38, Pea38c, Pea31c, Pea31d]. Studies [Wis38, Ter31]. Study [Bat33, Bat34, BD35, Cle37, Col33, HM36, Kar33, KM33, Le 31, Mah30, Mar36, MH31, Mor35, MCA36, Mün36, Nai39, Ris39, SM32, Woo30, WM34, Woo37, Die34]. Style [Yul39]. such [PS31]. Suffolk [Coc35]. Suggested [Pea35e]. Sums [Yul38b]. Supposed [Pea33c]. Surface [Ano30d, Cha32, Rie32]. Surfaces [Pre30]. Survey [Mor39a, VM38]. Sweden [Mah30]. Symmetrical [Gea33, Han34, PS31]. Systematic [Jef39, Yat39].

T [Rom33]. Tabellen [Har34]. Table [Dav34a, EM31, Hal38b, Hal39b]. Tables [Ano30d, Kar33, PS31, Har34]. Taken [Mor35, Pea31c, Til33]. Tasmanian [Mor39a]. Tasmanians [Wun39]. Tchebycheff [Ber38]. Tell [Ris39]. Term [Pea30a]. Terms [New39]. Test [Bis39, Bra33, Cha38, Dav34a, Dav39, Gea35b, Hal37, Hal39a, Law38, Law39, Mor39c, Pea32b, Pit38, Wel37]. Tested [Dav39, WT37]. Testing [Pea38b, PW35b]. Tests [GH37, Pea30b, Pea31b, Pea38b, Pit38, Pit39c, Wel38a, Mah35]. Their [Pea34a, Wis38]. Theorem [Ber38, Cam32, PW36]. Theoretical [RDM39, Pea33d]. Théorie [Die34]. Theory [Kon30, PS31, Pea39a, Ste31, Yul38a, Die34]. Third [HM36, Pep32]. Thirteenth [Col33]. Those [LY35]. Thoughts [Pea35e]. Three [BD35, Per33]. Times [KKS39]. Tippett [Nai38]. Tools [PS36]. Tooth [HM36]. Total [Cha32]. Transformation [Pea38b]. True [Dav39, WT37]. Tschebyschew [Ano30c]. Twelfth [Col33]. Twelve [Bay38]. Twenty [S.39]. Twenty-Five [S.39]. Two [Ber38, Cam32, Fin38, Hir37, Hot36, Mar32, Mor39c, Pea31c, Pea32c, PW33, Tho33, Wel35, Wel38b, Yul39, Mah35]. two-dimensional [Mah35]. Two-Factor [Cam32]. Type [Wic33]. Types [Mar32].

Ultimate [HM36]. Unbiased [Dav37]. Unequal [Wel38b]. United [VM38]. Unity [Dav33, Dav34b]. Univariate [Yul38b]. Universe [KKS39, McK33b, McK33a, McK35]. Unknown [Tho33]. Untersuchungen [Mah35, Ter31]. Unusual [Pea30a]. Use [CP34, Mor39b, Les33]. Used [Hal37, Hal39a]. Uses [Pea33b]. Usher [Pea35c]. Uspensky [Dav38].

V [Dav38]. Validity [Wiś35, Wis36]. Value [Coc38, Hal37, Hal39a, Pea33b, Pea33e]. Values [Dav34a, EM31]. Variables [Ber38, PW33, Ste34, Wel35]. Variance [Le 31, Pea31a, PP32, Pit38, Wil32]. Variances [Bis39, Fin38, Mor39c, Wel38b]. Variate [Ano30d, Fin38, Gar33]. Variates [Hot36, Pea31c, Pea33e]. Variation [Pea31a]. variety [Bat35].

Vaulted [SM32]. Verteilungen [Mah35]. VI [Pea35d]. View [Tho33]. Vision [TE35]. Vol [Pea35e]. Volume [Ano31e, Ano32g, Ano33h, Ano34h, Ano35e, Ano36g, Ano38e, Ano39e, THM^{+}32, Pea35c]. Volumes [Ano30d].

W [H.39, Lun38]. Wagner [Pea35d]. Want [EW32]. Welch [Pea35e]. Wellcome [Ris39]. West [You31]. where [Wil35]. Whether [Pea33c]. Which [Dav33, Dav34b, KKS39, Pea33e, Pit38]. Wilkinson [PM34]. William [Ano33e]. Wiśniewski’s [Pea35b]. Wissenschaft [Har34]. Woman [Les33]. Work [Pea36a, Pea38a, PW35b]. Wunderly [Mor39a].

x [Dav34b]. XXVII [Pea35e].

Years [BD35, S.39]. York [MS36].

z [Wel37]. z-Test [Wel37]. Zero [EM31]. zufallsartig [Mah35]. Zur [Von31]. Zwecke [Har34]. Zweidimensionale [Mah35].

References

Anderson:1933:EAS

[And33] Drysdale Anderson. An empirical age scale. Biometrika, 25(1/2): 61–70, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

          3510 (electronic). URL http://www.jstor.org/stable/2332261.

Anonymous:1930:BM

[Ano30a] Anonymous. Back matter. Biometrika, 22(1/2):??, July 1930. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

          URL http://www.jstor.org/stable/2332075.

Anonymous:1930:FM

[Ano30b] Anonymous. Front matter. Biometrika, 22(1/2):??, July 1930. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

          URL http://www.jstor.org/stable/2332065.

Anonymous:1930:IBP

[Ano30c] Anonymous. [illustration]: Biometrika portrait series: Tschebyschew, 1821–1894. Biometrika, 22(1/2):??, July 1930. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

          jstor.org/stable/2332066.

Anonymous:1930:RTD

[Ano30d] Anonymous. On the remaining tables for determining the volumes of a bi-variate normal surface. Biometrika, 22(1/2):1–35, July 1930.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332067.

Anonymous:1931:BMb

[Ano31a] Anonymous. Back matter. Biometrika, 22(3/4):??, May 1931. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332107.

Anonymous:1931:BMa

[Ano31b] Anonymous. Back matter. Biometrika, 23(1/2):??, November 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333634.

Anonymous:1931:FMb

[Ano31c] Anonymous. Front matter. Biometrika, 22(3/4):??, May 1931. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332094.

Anonymous:1931:FMa

[Ano31d] Anonymous. Front matter. Biometrika, 23(1/2):??, November 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333626.

Anonymous:1931:VI

[Ano31e] Anonymous. Volume information. Biometrika, 22(3/4):i–iv, May 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332093.

Anonymous:1932:BMa

[Ano32a] Anonymous. Back matter. Biometrika, 24(1/2):??, May 1932. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333805.

Anonymous:1932:BMb

[Ano32b] Anonymous. Back matter. Biometrika, 24(3/4):??, November 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331982.

Anonymous:1932:FMa

[Ano32c] Anonymous. Front matter. Biometrika, 24(1/2):??, May 1932. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333790.

Anonymous:1932:FMb

[Ano32d] Anonymous. Front matter. Biometrika, 24(3/4):??, November 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331969.

Anonymous:1932:IBP

[Ano32e] Anonymous. [illustration]: Biometrika portrait series: A. A. Markoff, 1856–1922. Biometrika, 24(3/4):??, November 1932. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331970.

Anonymous:1932:PRP

[Ano32f] Anonymous. Professor Rietz’s problem. Biometrika, 24(1/2):290– 291, May 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333803.

Anonymous:1932:VI

[Ano32g] Anonymous. Volume information. Biometrika, 24(3/4):i–v, November 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331968.

Anonymous:1933:BMa

[Ano33a] Anonymous. Back matter. Biometrika, 25(1/2):??, May 1933. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332281.

Anonymous:1933:BMb

[Ano33b] Anonymous. Back matter. Biometrika, 25(3/4):??, December 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332295.

Anonymous:1933:FMa

[Ano33c] Anonymous. Front matter. Biometrika, 25(1/2):??, May 1933. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332257.

Anonymous:1933:FMb

[Ano33d] Anonymous. Front matter. Biometrika, 25(3/4):??, December 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332283.

Anonymous:1933:IBP

[Ano33e] Anonymous. [illustration]: Biometrika portrait series: William Letty Kn. Biometrika, 25(3/4):??, December 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332294.

Anonymous:1933:NMM

[Ano33f] Anonymous. Note on Mr McKay’s paper. Biometrika, 25(1/2):210– 213, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332279.

Anonymous:1933:NFF

[Ano33g] Anonymous. Note on the fitting of frequency curves. Biometrika, 25(1/2):213–216, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332280.

Anonymous:1933:VI

[Ano33h] Anonymous. Volume information. Biometrika, 25(3/4):i–viii, December 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332282.

Anonymous:1934:BMa

[Ano34a] Anonymous. Back matter. Biometrika, 26(1/2):??, May 1934. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332064.

Anonymous:1934:BMb

[Ano34b] Anonymous. Back matter. Biometrika, 26(3):??, December 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332203.

Anonymous:1934:BMc

[Ano34c] Anonymous. Back matter. Biometrika, 26(4):??, December 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331991.

Anonymous:1934:FMa

[Ano34d] Anonymous. Front matter. Biometrika, 26(1/2):??, May 1934. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332051.

Anonymous:1934:FMb

[Ano34e] Anonymous. Front matter. Biometrika, 26(3):??, December 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332201.

Anonymous:1934:FMc

[Ano34f] Anonymous. Front matter. Biometrika, 26(4):??, December 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331984.

Anonymous:1934:RPS

[Ano34g] Anonymous. Remarks on Professor Steffensen’s measure of contingency. Biometrika, 26(1/2):255–260, May 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332059.

Anonymous:1934:VI

[Ano34h] Anonymous. Volume information. Biometrika, 26(4):i–xii, December 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331983.

Anonymous:1935:BMa

[Ano35a] Anonymous. Back matter. Biometrika, 27(1/2):??, March 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332050.

Anonymous:1935:BMb

[Ano35b] Anonymous. Back matter. Biometrika, 27(3/4):??, October 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332704.

Anonymous:1935:FMa

[Ano35c] Anonymous. Front matter. Biometrika, 27(1/2):??, March 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332036.

Anonymous:1935:FMb

[Ano35d] Anonymous. Front matter. Biometrika, 27(3/4):??, October 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332690.

Anonymous:1935:VI

[Ano35e] Anonymous. Volume information. Biometrika, 27(3/4):i–vii, October 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332689.

Anonymous:1936:BMa

[Ano36a] Anonymous. Back matter. Biometrika, 28(1/2):??, June 1936. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334133.

Anonymous:1936:BMb

[Ano36b] Anonymous. Back matter. Biometrika, 28(3/4):??, December 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333961.

Anonymous:1936:FMa

[Ano36c] Anonymous. Front matter. Biometrika, 28(1/2):??, June 1936. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334119.

Anonymous:1936:FMb

[Ano36d] Anonymous. Front matter. Biometrika, 28(3/4):??, December 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333949.

Anonymous:1936:I

[Ano36e] Anonymous. [illustration]. Biometrika, 28(3/4):??, December 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333950.

Anonymous:1936:KP

[Ano36f] Anonymous. Karl Pearson, 1857–1936. Biometrika, 28(1/2):??, June 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334120.

Anonymous:1936:VI

[Ano36g] Anonymous. Volume information. Biometrika, 28(3/4):i–vii, December 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333948.

Anonymous:1937:BM

[Ano37a] Anonymous. Back matter. Biometrika, 29(1/2):??, June 1937. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332416.

Anonymous:1937:FM

[Ano37b] Anonymous. Front matter. Biometrika, 29(1/2):??, June 1937. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332403.

Anonymous:1938:BMb

[Ano38a] Anonymous. Back matter. Biometrika, 29(3/4):??, February 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332017.

Anonymous:1938:BMa

[Ano38b] Anonymous. Back matter. Biometrika, 30(1/2):??, June 1938. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332240.

Anonymous:1938:FMb

[Ano38c] Anonymous. Front matter. Biometrika, 29(3/4):??, February 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332004.

Anonymous:1938:FMa

[Ano38d] Anonymous. Front matter. Biometrika, 30(1/2):??, June 1938. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332218.

Anonymous:1938:VI

[Ano38e] Anonymous. Volume information. Biometrika, 29(3/4):i–vii, February 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332003.

Anonymous:1939:BMb

[Ano39a] Anonymous. Back matter. Biometrika, 30(3/4):??, January 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332662.

Anonymous:1939:BMa

[Ano39b] Anonymous. Back matter. Biometrika, 31(1/2):??, July 1939. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334989.

Anonymous:1939:FMb

[Ano39c] Anonymous. Front matter. Biometrika, 30(3/4):??, January 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332646.

Anonymous:1939:FMa

[Ano39d] Anonymous. Front matter. Biometrika, 31(1/2):??, July 1939. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334969.

Anonymous:1939:VI

[Ano39e] Anonymous. Volume information. Biometrika, 30(3/4):i–viii, January 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332645.

Anonymous:1948:CMR

[Ano48a] Anonymous. Corrigenda: “Moments of the Ratio of the Mean Deviation to the Standard Deviation for Normal Samples”. Biometrika, 35(3/4):424, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332365. See [Gea36].

Anonymous:1948:CPM

[Ano48b] Anonymous. Corrigenda: “The Probability of the Mean Deviation”. Biometrika, 35(3/4):424, December 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332366. See [Gea36, Pea45].

Ayyangar:1934:NIM

[Ayy34a] A. A. Krishnaswami Ayyangar. A note on the incomplete moments of the hypergeometrical series. Biometrika, 26(1/2):264–265, May 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332062.

Ayyangar:1934:NRF

[Ayy34b] A. A. Krishnaswami Ayyangar. Note on the recurrence formulae for the moments of the point binomial. Biometrika, 26(1/2):262–264, May 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332061.

B:1936:BRB

[B.36] J. B. Book review: J. Arthur Harris, Botanist and Biometrician by C. Otto Rosendahl; Ross Aiken Gortner; George O. Burr. Biometrika, 28(3/4):444, December 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333960.

Baten:1933:SSB

[Bat33] William Dowell Baten. A statistical study of the Daucus Carota L. Biometrika, 25(1/2):186–195, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332274.

Baten:1934:SSB

[Bat34] William Dowell Baten. A statistical study of the Daucus Carota L. (second article). Biometrika, 26(4):443–468, December 1934. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331989.

Baten:1935:CNL

[Bat35] William Dowell Baten. Constancy in the number of ligulate flowers of Chrysanthemum Leucanthemium, variety Pinnatifidum, during the flowering season. Biometrika, 27(1/2):260–266, March 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332046.

Baten:1936:IPS

[Bat36] William Dowell Baten. Influence of position on structure of inflorescences of Cicuta Maculata. Biometrika, 28(1/2):64–83, June 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334125.

Bayly:1938:GQF

[Bay38] B. De F. Bayly. Gauss’ quadratic formula with twelve ordinates. Biometrika, 30(1/2):193–194, June 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332235.

Bayley:1935:GCB

[BD35] Nancy Bayley and Frank C. Davis. Growth changes in bodily size and proportions during the first three years: A developmental study of sixty-one children by repeated measurements. Biometrika, 27(1/2):26–87, March 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332039.

Beall:1939:MEP

[Bea39] Geoffrey Beall. Methods of estimating the population of insects in a field. Biometrika, 30(3/4):422–439, January 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332657.

Bell:1933:PBN

[Bel33] Julia Bell. Plural births with a new pedigree. Biometrika, 25 (1/2):110–120, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332266.

Berge:1938:NFT

[Ber38] P. O. Berge. A note on a form of Tchebycheff’s theorem for two variables. Biometrika, 29(3/4):405–406, February 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332015.

Bishop:1939:CTH

[Bis39] D. J. Bishop. On a comprehensive test for the homogeneity of variances and covariances in multivariate problems. Biometrika, 31(1/2):31–55, July 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334974.

Brandner:1933:TSD

[Bra33] Fred A. Brandner. A test of the significance of the difference of the correlation coefficients in normal bivariate samples. Biometrika, 25(1/2):102–109, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332265.

Brash:1935:BAS

[BY35] J. C. Brash and Matthew Young. The Bidford-on-Avon skulls. Biometrika, 27(3/4):373–387, October 1935. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332698.

Camp:1932:CST

[Cam32] Burton H. Camp. The converse of Spearman’s two-factor theorem. Biometrika, 24(3/4):418–427, November 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331975.

Camp:1934:SGF

[Cam34] Burton H. Camp. Spearman’s general factor again. Biometrika, 26(1/2):260–261, May 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332060.

Chapelin:1932:MPP

[Cha32] Jacques Chapelin. On a method of proceeding from partial cell frequencies to ordinates and to total cell frequencies in the case of a bivariate frequency surface. Biometrika, 24(3/4):495–497, November 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2331980.

Chapman:1938:ATP

[Cha38] R. A. Chapman. Applicability of the z test to a Poisson distribution. Biometrika, 30(1/2):188–190, June 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332233.

Church:1932:NMR

[Chu32] A. E. R. Church. Note on a memoir by A. E. R. Church. Biometrika, 24(1/2):292, May 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333804.

Cleaver:1937:CBS

[Cle37] Frank H. Cleaver. A contribution to the biometric study of the human mandible. Biometrika, 29(1/2):80–112, June 1937. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332408.

Cheeseman:1939:DE

[CMR39] E. A. Cheeseman, W. J. Martin, and W. T. Russell. Disease and environment. Biometrika, 30(3/4):341–362, January 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332654.

Cockayne:1935:SPF

[Coc35] E. A. Cockayne. A second piebald family from Suffolk. Biometrika, 27(1/2):1–4, March 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332037.

Cockayne:1936:CPH

[Coc36] E. A. Cockayne. Cleft palate, hare lip, dacryocystitis, and cleft hand and feet. Biometrika, 28(1/2):60–63, June 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334124.

Cochran:1938:NJB

[Coc38] W. G. Cochran. Note on J. B. S. Haldane’s paper: “The Exact Value of the Moments of the Distribution of χ^{2}”. Biometrika, 29(3/4):407, February 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332016.

Collett:1933:STT

[Col33] Margot Collett. A study of twelfth and thirteenth dynasty skulls from Kerma (Nubia). Biometrika, 25(3/4):254–284, December 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332285.

Clopper:1934:UCF

[CP34] C. J. Clopper and E. S. Pearson. The use of confidence or fiducial limits illustrated in the case of the binomial. Biometrika, 26(4):404–413, December 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331986.

Davies:1932:BQD

[Dav32] Owen L. Davies. On the betas of quadrilateral distributions. Biometrika, 24(3/4):498–510, November 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331981.

Davies:1933:AFH

[Dav33] O. L. Davies. On asymptotic formulae for the hypergeometric series: I. Hypergeometric series in which the fourth element, x, is unity.

Biometrika, 25(3/4):295–322, December 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332287.

David:1934:TRR

[Dav34a] Florence N. David. On the P_{λ_{n}} test for randomness: Remarks, further illustration, and table of P_{λ_{n}} for given values of − log_{10} λ_{n}. Biometrika, 26(1/2):1–11, May 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332052.

Davies:1934:AFH

[Dav34b] O. L. Davies. On asympotic formulae for the hypergeometric series: Hypergeometric series in which the fourth element, x, is not necessarily unity. Biometrika, 26(1/2):59–107, May 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332054.

David:1937:NUL

[Dav37] F. N. David. A note on unbiased limits for the correlation coefficient. Biometrika, 29(1/2):157–160, June 1937. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332415.

David:1938:BRB

[Dav38] F. N. David. Book review: Introduction to Mathematical Probability by J. V. Uspensky. Biometrika, 30(1/2):194–195, June 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332236.

David:1939:NST

[Dav39] F. N. David. On Neyman’s “smooth” test for goodness of fit: I. Distribution of the criterion ψ^{2} when the hypothesis tested is true. Biometrika, 31(1/2):191–199, July 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334982.

Dieulefait:1934:CET

[Die34] Carlos E. Dieulefait. Contribution à l’étude de la théorie de la correlation. (French) [Contribution to the study of the theory of correlation]. Biometrika, 26(4):379–403, December 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331985.

Dingwall:1933:SED

[DY33] Doris Dingwall and Matthew Young. The skulls from excavations at Dunstable, Bedfordshire. Biometrika, 25(1/2):147–157, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332269.

ElShanawany:1936:IAA

[El 36] M. R. El Shanawany. An illustration of the accuracy of the χ^{2} approximation. Biometrika, 28(1/2):179–187, June 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334129.

Elderton:1933:AMJ

[Eld33] W. Palin Elderton. Adjustments for the moments of J-shaped curves. Biometrika, 25(1/2):179–180, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332271. See note [Pea33a].

Elderton:1931:TVD

[EM31] Ethel M. Elderton and Margaret Moul. Table of the values of the differences of the powers of zero. Biometrika, 22(3/4):306–308, May 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332098.

Elderton:1932:NWN

[EW32] E. M. Elderton and T. L. Woo. On the normality or want of normality in the frequency distributions of cranial measurements. Biometrika, 24(1/2):45–54, May 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333795.

Fieller:1931:DP

[Fie31a] E. C. Fieller. The duration of play. Biometrika, 22(3/4):377–404, May 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332102.

Fieller:1931:PP

[Fie31b] E. C. Fieller. A problem in probability. Biometrika, 22(3/4):425– 427, May 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332106.

Fieller:1932:DIN

[Fie32] E. C. Fieller. The distribution of the index in a normal bivariate population. Biometrika, 24(3/4):428–440, November 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331976.

Finney:1938:DRE

[Fin38] D. J. Finney. The distribution of the ratio of estimates of the two variances in a sample from a normal bi-variate population. Biometrika, 30(1/2):190–192, June 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332234.

Garwood:1933:PIC

[Gar33] F. Garwood. The probability integral of the correlation coefficient in samples from a normal bi-variate population. Biometrika, 25(1/2): 71–78, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332262.

Garwood:1936:FLP

[Gar36] F. Garwood. Fiducial limits for the Poisson distribution. Biometrika, 28(3/4):437–442, December 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333958.

Geary:1933:GEM

[Gea33] R. C. Geary. A general expression for the moments of certain symmetrical functions of normal samples. Biometrika, 25(1/2):184–186, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332273.

Geary:1935:NCB

[Gea35a] R. C. Geary. Note on the correlation between β_{2} and w′. Biometrika, 27(3/4):353–355, October 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332695.

Geary:1935:RMD

[Gea35b] R. C. Geary. The ratio of the mean deviation to the standard deviation as a test of normality. Biometrika, 27(3/4):310–332, October 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332693.

Geary:1936:MRM

[Gea36] R. C. Geary. Moments of the ratio of the mean deviation to the standard deviation for normal samples. Biometrika, 28(3/4):295–307, December 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333953.

See corrigenda [Ano48a] and comments [SBL79].

Gruneberg:1937:TGF

[GH37] Hans Grüneberg and J. B. S. Haldane. Tests of goodness of fit applied to records of Mendelian segregation in mice. Biometrika, 29(1/2):144–153, June 1937. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332412.

Gordon:1939:EBP

[Gor39] Robert Dean Gordon. Estimating bacterial populations by the dilution method. Biometrika, 31(1/2):167–180, July 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334979.

Greenwood:1938:BDN

[GR38] Major Greenwood and W. T. Russell. Bright’s disease, nephritis and arterio-sclerosis: A contribution to the history of medical statistics. Biometrika, 29(3/4):249–276, February 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332006.

H:1939:BRB

[H.39] J. B. S. H. Book review: Principles of Genetics by E. W. Sinnott; L. C. Dunn. Biometrika, 31(1/2):220, July 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334986.

Haldane:1937:EVM

[Hal37] J. B. S. Haldane. The exact value of the moments of the distribution of χ^{2}, used as a test of goodness of fit, when expectations are small. Biometrika, 29(1/2):133–143, June 1937. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor. org/stable/2332411. See correction [Hal39a].

Haldane:1938:ANC

[Hal38a] J. B. S. Haldane. The approximate normalization of a class of frequency distributions. Biometrika, 29(3/4):392–404, February 1938.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332014.

Haldane:1938:FSM

[Hal38b] J. B. S. Haldane. The first six moments of χ^{2} for an n-fold table with n degrees of freedom when some expectations are small. Biometrika, 29(3/4):389–391, February 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2332013. See correction [Hal39b].

Haldane:1938:HNP

[Hal38c] J. B. S. Haldane. Heterostylism in natural populations of the primrose, Primula Acaulis. Biometrika, 30(1/2):196–198, June 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332237.

Haldane:1939:CEV

[Hal39a] J. B. S. Haldane. Correction: “The Exact Value of the Moments of the Distribution of χ^{2}, Used as a Test of Goodness of Fit, When Expectations are Small”. Biometrika, 31(1/2):220, July 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334987. See [Hal37].

Haldane:1939:CFS

[Hal39b] J. B. S. Haldane. Correction: “The First Six Moments of χ^{2} for an n-Fold Table with n Degrees of Freedom when some Expectations are Small”. Biometrika, 31(1/2):220, July 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334988. See [Hal38b].

Haldane:1939:NPA

[Hal39c] J. B. S. Haldane. Note on the preceding analysis of Mendelian segregations. Biometrika, 31(1/2):67–71, July 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334976.

Hansmann:1934:CNN

[Han34] G. H. Hansmann. On certain non-normal symmetrical frequency distributions. Biometrika, 26(1/2):129–195, May 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332056.

Harzer:1934:TAS

[Har34] Paul Harzer. Tabellen fur alle Statistische Zwecke in Wissenschaft und Praxis (German) [Statistical tables for all purposes in science and practice]. Biometrika, 26(4):469–470, December 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331990.

Hendricks:1937:AMM

[Hen37] Walter A. Hendricks. An application of the method of maximum likelihood. Biometrika, 29(1/2):154–155, June 1937. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332413.

Hey:1938:NME

[Hey38] G. B. Hey. A new method of experimental sampling illustrated on certain non-normal populations. Biometrika, 30(1/2):68–80, June 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332225.

Hirschfeld:1937:DRC

[Hir37] H. O. Hirschfeld. The distribution of the ratio of covariance estimates in two samples drawn from normal bivariate populations. Biometrika, 29(1/2):65–79, June 1937. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332407.

Henry:1936:PSE

[HM36] C. Bowdler Henry and G. M. Morant. A preliminary study of the eruption of the mandibular third molar tooth in man based on measurements obtained from radiographs, with special reference to the problem of predicting cases of ultimate impaction of the tooth. Biometrika, 28(3/4):378–427, December 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333956.

Hojo:1933:FNR

[Hoj33] Tokishige Hojo. A further note on the relation between the median and the quartiles in small samples from a normal population. Biometrika, 25(1/2):79–90, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332263.

Hotelling:1936:RBT

[Hot36] Harold Hotelling. Relations between two sets of variates. Biometrika, 28(3/4):321–377, December 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333955.

Jeffreys:1939:RSA

[Jef39] Harold Jeffreys. Random and systematic arrangements. Biometrika, 31(1/2):1–8, July 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334970.

Johnson:1939:CCD

[JW39] N. L. Johnson and B. L. Welch. On the calculation of the cumulants of the χ-distribution. Biometrika, 31(1/2):216–218, July 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334984.

Karn:1933:FSM

[Kar33] M. Noel Karn. A further study of methods of constructing life tables when certain causes of death are eliminated. Biometrika, 25(1/2):91–101, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332264.

Kendall:1938:NMR

[Ken38] M. G. Kendall. A new measure of rank correlation. Biometrika, 30(1/2):81–93, June 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332226.

Kendall:1939:DSC

[KKS39] M. G. Kendall, Sheila F. H. Kendall, and B. Babington Smith. The distribution of Spearman’s coefficient of rank correlation in a universe in which all rankings occur an equal number of times. Biometrika, 30(3/4):251–273, January 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332649.

Kitson:1933:SNS

[KM33] Elisabeth Kitson and G. M. Morant. A study of the Naga skull. Biometrika, 25(1/2):1–20, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332258.

Kolodziejczyk:1935:ICS

[Kol35] Stanislaw Kolodziejczyk. On an important class of statistical hypotheses. Biometrika, 27(1/2):161–190, March 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332043.

Kondo:1930:TSD

[Kon30] T. Kondo. A theory of the sampling distribution of standard deviations. Biometrika, 22(1/2):36–64, July 1930. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332068.

Lawley:1938:GFT

[Law38] D. N. Lawley. A generalization of Fisher’s z test. Biometrika, 30(1/2):180–187, June 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2332232. See correction [Law39].

Lawley:1939:CGF

[Law39] D. N. Lawley. A correction to “A Generalization of Fisher’s z Test”. Biometrika, 30(3/4):467–469, January 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332659. See [Law38].

LeRoux:1931:SDV

[Le 31] J. M. Le Roux. A study of the distribution of the variance in small samples. Biometrika, 23(1/2):134–190, November 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333632.

Ledermann:1939:SDS

[Led39] Walter Ledermann. Sampling distribution and selection in a normal population. Biometrika, 30(3/4):295–304, January 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332651.

Lesser:1933:NSP

[Les33] Pamela C. V. Lesser. Note on the shrinkage of physical characters in man and woman with age, as an illustration of the use of χ^{2}, P methods. Biometrika, 25(1/2):197–202, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332276.

Luders:1934:SSE

[Lud34] Rolf Luders. Die Statistik der Seltenen Ereignisse. (German) [The statistics of rare events]. Biometrika, 26(1/2):108–128, May 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332055.

Lundberg:1938:BRB

[Lun38] O. Lundberg. Book review: Frequency Curves and Correlation by W. Palin Elderton. Biometrika, 30(1/2):203–204, June 1938. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332239.

Layard:1935:BSI

[LY35] Doris Layard and Matthew Young. The Burwell skulls, including a comparison with those of certain other Anglo–Saxon series. Biometrika, 27(3/4):388–408, October 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332699.

Mahalanobis:1930:SSC

[Mah30] P. C. Mahalanobis. A statistical study of certain anthropometric measurements from Sweden. Biometrika, 22(1/2):94–108, July 1930. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332070.

Mahlmann:1935:BUZ

[Mah35] Henry Mahlmann. Ein Beitrag zu Untersuchungen über Zweidimensionale Verteilungen von Massenpunkten bei zufallsartig bedingten Bewegungen. (German) [A contribution to tests on two-dimensional distributions of mass points in conditional random-like movements]. Biometrika, 27(1/2):191–226, March 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332044.

Martin:1932:CAT

[Mar32] E. S. Martin. A comparison of the accuracy of two types of quadrature formulae. Biometrika, 24(1/2):280–287, May 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333801.

Martin:1934:CMC

[Mar34] E. S. Martin. On corrections for the moment coefficients of frequency distributions when the start of the frequency is one of the characteristics to be determined. Biometrika, 26(1/2):12–58, May 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332053.

Martin:1936:SES

[Mar36] E. S. Martin. A study of an Egyptian series of mandibles, with special reference to mathematical methods of sexing. Biometrika, 28(1/2):149–178, June 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334128.

Martin:1938:RMC

[MC38] W. J. Martin and E. A. Cheeseman. Rural mortality. Its comparative sex incidence. Biometrika, 30(1/2):57–64, June 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332223.

Morant:1936:BSH

[MCA36] G. M. Morant, Margot Collett, and N. K. Adyanthaya. A biometric study of the human mandible. Biometrika, 28(1/2):84–122, June 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334126.

McKay:1932:BFD

[McK32] A. T. McKay. A Bessel function distribution. Biometrika, 24(1/2): 39–44, May 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333794.

McKay:1933:DSN

[McK33a] A. T. McKay. The distribution of β_{2} in samples of 4 from a normal universe. Biometrika, 25(3/4):411–415, December 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332291.

McKay:1933:DSF

[McK33b] A. T. McKay. The distribution of \sqrt{beta_{1}} in samples of four from a normal universe. Biometrika, 25(1/2):204–210, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332278.

McKay:1935:DDB

[McK35] A. T. McKay. The distribution of the difference between the extreme observation and the sample mean in samples of n from a normal universe. Biometrika, 27(3/4):466–471, October 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332702.

McMullen:1939:SM

[McM39] L. McMullen. “Student” as a man. Biometrika, 30(3/4):205–210, January 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332647.

Morant:1931:SRE

[MH31] G. M. Morant and M. F. Hoadley. A study of the recently excavated Spitalfields crania. Biometrika, 23(1/2):191–248, November 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333633.

Moore:1935:HMH

[Moo35] Eldon Moore. Heredity, mainly human. Biometrika, 27(3/4):471, October 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332703.

Morant:1935:SPE

[Mor35] G. M. Morant. A study of predynastic Egyptian skulls from Badari based on measurements taken by Miss B. N. Stoessiger and Professor D. E. Derry. Biometrika, 27(3/4):293–309, October 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332692.

Morant:1937:CEC

[Mor37] G. M. Morant. A contribution to Eskimo craniology based on previously published measurements. Biometrika, 29(1/2):1–20, June 1937. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332404.

Morant:1938:DNH

[Mor38] G. M. Morant. A description of nine human skulls from Iran excavated by Sir Aurel Stein, K.C.I.E. Biometrika, 30(1/2):130–133, June 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332228.

Morant:1939:NDJ

[Mor39a] G. M. Morant. Note on Dr. J. Wunderly’s survey of Tasmanian crania. Biometrika, 30(3/4):338–340, January 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332653.

Morant:1939:USM

[Mor39b] G. M. Morant. The use of statistical methods in the investigation of problems of classification in anthropology: Part I. The general nature of the material and the form of intraracial distributions of metrical characters. Biometrika, 31(1/2):72–98, July 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334977.

Morgan:1939:TSD

[Mor39c] W. A. Morgan. A test for the significance of the difference between the two variances in a sample from a normal bivariate population. Biometrika, 31(1/2):13–19, July 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334972.

McKay:1933:NDR

[MP33] A. T. McKay and E. S. Pearson. A note on the distribution of range in samples of n. Biometrika, 25(3/4):415–420, December 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332292.

Morant:1936:EID

[MS36] G. M. Morant and Otto Samson. An examination of investigations by Dr Maurice Fishberg and Professor Franz Boas dealing with measurements of Jews in New York. Biometrika, 28(1/2):1–31, June 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334121.

Muller:1931:ACF

[Mul31] J. H. Muller. On the application of continued fractions to the evaluation of certain integrals, with special reference to the incomplete beta-function. Biometrika, 22(3/4):284–297, May 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332096.

Munter:1936:SLL

[Mün36] Phil. A. Heinrich Münter. A study of the lengths of the long bones of the arms and legs in man, with special reference to Anglo–Saxon skeletons. Biometrika, 28(3/4):258–294, December 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333952.

Nair:1936:NKP

[Nai36a] K. Raghavan Nair. Note on Karl Pearson’s paper: “On a Method of Ascertaining Limits to the Actual Number of Marked Members in a Population of Given Size from a Sample”. Biometrika, 28 (3/4):442–443, December 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333959.

Nair:1936:SEG

[Nai36b] U. S. Nair. The standard error of Gini’s mean difference. Biometrika, 28(3/4):428–436, December 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333957.

Nair:1938:TRS

[Nai38] K. N. Nair. On Tippett’s random sampling numbers. Sankhyā (Indian Journal of Statistics), Series A. Methods and Techniques, 4(1): 65–72, 1938. CODEN SNKYA5. ISSN 0036-4452. URL http://

www.jstor.org/stable/40383888.

Nair:1939:AMF

[Nai39] U. S. Nair. The application of the moment function in the study of distribution laws in statistics. Biometrika, 30(3/4):274–294, January 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332650.

Newman:1939:DRS

[New39] D. Newman. The distribution of range in samples from a normal population, expressed in terms of an independent estimate of standard deviation. Biometrika, 31(1/2):20–30, July 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334973.

Neyman:1938:HNK

[Ney38] J. Neyman. A historical note on Karl Pearson’s deducation of the moments of the binomial. Biometrika, 30(1/2):11–15, June 1938.

CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332220.

Neyman:1931:FND

[NP31] J. Neyman and E. S. Pearson. Further notes on the χ^{2} distribution. Biometrika, 22(3/4):298–305, May 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332097.

Neyman:1938:NSP

[NP38] J. Neyman and E. S. Pearson. Note on some points in “Student’s” Paper on “Comparison Between Balanced and Random Arrangements of Field Plots”. Biometrika, 29(3/4):379–388, February 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332012.

Nussey:1938:PF

[Nus38] A. M. Nussey. A piebald family. Biometrika, 30(1/2):65–67, June 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332224.

Pearce:1930:UFD

[Pea30a] Thos. Vibert Pearce. An unusual frequency distribution — the term of abortion. Biometrika, 22(1/2):250–252, July 1930. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332074.

Pearson:1930:FDT

[Pea30b] E. S. Pearson. A further development of tests for normality. Biometrika, 22(1/2):239–249, July 1930. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332073.

Pearson:1931:AVC

[Pea31a] Egon S. Pearson. The analysis of variance in cases of non-normal variation. Biometrika, 23(1/2):114–133, November 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333631.

Pearson:1931:NTN

[Pea31b] Egon S. Pearson. Note on tests for normality. Biometrika, 22 (3/4):423–424, May 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332104.

Pearson:1931:NRB

[Pea31c] Karl Pearson. On the nature of the relationship between two of “Student’s” variates (Z_{1} and Z_{2}) when samples are taken from a bivariate normal population. Biometrika, 22(3/4):405–422, May 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332103.

Pearson:1931:SPS

[Pea31d] Karl Pearson. Some properties of “Student’s” z: Correlation, regression and scedasticit of z with the mean and standard deviation of the sample. Biometrika, 23(1/2):1–9, November 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333627.

Pearson:1932:PLD

[Pea32a] Egon S. Pearson. The percentage limits for the distribution of range in samples from a normal population (n < 100). Biometrika, 24 (3/4):404–417, November 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331974.

Pearson:1932:EDT

[Pea32b] Karl Pearson. Experimental discussion of the (χ^{2}, P) test for goodness of fit. Biometrika, 24(3/4):351–381, November 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331972.

Pearson:1932:PTI

[Pea32c] Karl Pearson. On the probability that two independent distributions of frequency are really samples from the same parent population. Biometrika, 24(3/4):457–470, November 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331978.

Pearson:1933:NMP

[Pea33a] K. Pearson. Note on Mr Palin Elderton’s corrections to the moments of J-curves. Biometrika, 25(1/2):180–184, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332272. See [Eld33].

Pearson:1933:CCS

[Pea33b] Karl Pearson. The cranial coordinatograph, the standard planes of the skull, and the value of Cartesian geometry to the craniologist, with some illustrations of the uses of the new method. Biometrika, 25(3/4):217–253, December 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332284.

Pearson:1933:MDW

[Pea33c] Karl Pearson. On a method of determining whether a sample of size n supposed to have been drawn from a parent population having a known probability integral has probably been drawn at random. Biometrika, 25(3/4):379–410, December 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332290.

Pearson:1933:ADB

[Pea33d] Karl Pearson. On the applications of the double Bessel function K_{τ_{1},τ_{2}}(X) to statistical problems: Part I. Theoretical. Biometrika, 25(1/2):158–178, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332270.

Pearson:1933:PPI

[Pea33e] Karl Pearson. On the parent population with independent variates which gives the minimum value of ϕ^{2} for a given sample. Biometrika, 25(1/2):134–146, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332268.

Pearson:1934:STH

[Pea34a] K. Pearson. On simometers and their handling. Biometrika, 26(1/2):265–268, May 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332063.

Pearson:1934:NMD

[Pea34b] Karl Pearson. On a new method of determining “goodness of fit”. Biometrika, 26(4):425–442, December 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www. jstor.org/stable/2331988. According to [Tei65, page 36], this paper introduced the P_{n}(λ) test for “determining whether a sample of size n, supposed to have been drawn at random from a parent population having a known probability integral has probably been drawn at random.” See [Nai38] for its first use on random numbers from a uniform distribution.

Pearson:1935:CMG

[Pea35a] E. S. Pearson. A comparison of β_{2} and Mr Geary’s w_{n} criteria. Biometrika, 27(3/4):333–352, October 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332694.

Pearson:1935:CBC

[Pea35b] K. Pearson. On the corrections for broad categories, being a note on Mr Wiśniewski’s memoir. Biometrika, 27(3/4):364–372, October 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332697. See

[Wiś35].

Pearson:1935:NDU

[Pea35c] Karl Pearson. Note on Dr Usher’s paper on epicanthus (pp. 5–25 of this volume). Biometrika, 27(1/2):271–272, March 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332049.

Pearson:1935:NSV

[Pea35d] Karl Pearson. Note on Section VI of Dr K. Wagner’s memoir. Biometrika, 27(1/2):133–144, March 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332041.

Pearson:1935:TSP

[Pea35e] Karl Pearson. Thoughts suggested by the papers of Messrs Welch and Kolodziejcyk (Biometrika, Vol. XXVII). Biometrika, 27(1/2):145–190, March 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332045.

Pearson:1936:KPA

[Pea36a] E. S. Pearson. Karl Pearson: An appreciation of some aspects of his life and work: Part I: 1857–1906. Biometrika, 28(3/4):193–257, December 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2333951.

Pearson:1936:MMM

[Pea36b] Karl Pearson. Method of moments and method of maximum likelihood. Biometrika, 28(1/2):34–59, June 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2334123.

Pearson:1936:JGR

[Pea36c] Karl Pearson. On Jewish–Gentile relationships. Biometrika, 28 (1/2):32–33, June 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334122.

Pearson:1937:MLM

[Pea37a] E. S. Pearson. Maximum likelihood and methods of estimation. Biometrika, 29(1/2):155–156, June 1937. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332414.

Pearson:1937:SAP

[Pea37b] E. S. Pearson. Some aspects of the problem of randomization. Biometrika, 29(1/2):53–64, June 1937. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332406.

Pearson:1938:ASA

[Pea38a] E. S. Pearson. An appreciation of some aspects of his life and work. Biometrika, 29(3/4):161–248, February 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332005.

Pearson:1938:PIT

[Pea38b] E. S. Pearson. The probability integral transformation for testing goodness of fit and combining independent tests of significance. Biometrika, 30(1/2):134–148, June 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332229.

Pearson:1938:SAP

[Pea38c] E. S. Pearson. Some aspects of the problem of randomization: II. An illustration of “Students” inquiry into the effect of “balancing” in agricultural experiment. Biometrika, 30(1/2):159–179, June 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332231.

Pearson:1939:NPP

[Pea39a] E. S. Pearson. Note on Professor Pitman’s contribution to the theory of estimation. Biometrika, 30(3/4):471–474, January 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332661.

Pearson:1939:NID

[Pea39b] E. S. Pearson. Note on the inverse and direct methods of estimation in R. D. Gordon’s problem. Biometrika, 31(1/2):181–186, July 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2334980.

Pearson:1939:SS

[Pea39c] E. S. Pearson. “Student” as statistician. Biometrika, 30(3/4): 210–250, January 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332648.

Pearson:1945:EN

[Pea45] E. S. Pearson. Editorial note. Biometrika, 33(3):252–253, November 1945. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332304. See

corrigenda [Ano48b].

Pepper:1932:SDT

[Pep32] Joseph Pepper. The sampling distribution of the third moment coefficient — an experiment. Biometrika, 24(1/2):55–64, May 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2333796.

Perlo:1933:DSR

[Per33] Victor Perlo. On the distribution of Student’s ratio for samples of three drawn from a rectangular distribution. Biometrika, 25(1/2):203–204, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332277.

Pertia:1934:RS

[Per34] Marioara Pertia. The roumanian silhouette. Biometrika, 26 (4):414–424, December 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331987.

Pitman:1938:STW

[Pit38] E. J. G. Pitman. Significance tests which May be applied to samples from any populations: III. The analysis of variance test. Biometrika, 29(3/4):322–335, February 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332008.

Pitman:1939:ELS

[Pit39a] E. J. G. Pitman. The estimation of the location and scale parameters of a continuous population of any given form. Biometrika, 30(3/4):391–421, January 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332656.

Pitman:1939:NNC

[Pit39b] E. J. G. Pitman. A note on normal correlation. Biometrika, 31(1/2): 9–12, July 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334971.

Pitman:1939:THC

[Pit39c] E. J. G. Pitman. Tests of hypotheses concerning location and scale parameters. Biometrika, 31(1/2):200–215, July 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334983.

Pearson:1934:WHO

[PM34] Karl Pearson and G. M. Morant. The Wilkinson head of Oliver Cromwell and its relationship to busts, masks and painted portraits. Biometrika, 26(3):1–116, December 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332202.

Pearson:1931:RDP

[PP31] Margaret Pearson and Karl Pearson. On the relation of the duration of pregnancy to size of litter and other characters in bitches. Biometrika, 22(3/4):309–323, May 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332099.

Pearson:1932:MCV

[PP32] Karl Pearson and Margaret V. Pearson. On the mean character and variance of a ranked individual, and on the mean and variance of the intervals between ranked individuals: Part II. Case of certain skew curves. Biometrika, 24(1/2):203–279, May 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333800.

Pearson:1935:NEH

[PP35] Karl Pearson and Margaret V. Pearson. On the numerical evaluation of high order incomplete Eulerian integrals. Biometrika, 27(3/4):409–423, October 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332700.

Pretorius:1930:SBF

[Pre30] S. J. Pretorius. Skew bivariate frequency surfaces, examined in the light of numerical illustrations. Biometrika, 22(1/2):109–223, July 1930. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332071.

Pearson:1931:TPI

[PS31] Karl Pearson and Brenda Stoessiger. Tables of the probability integrals of symmetrical frequency curves in the case of low powers such as arise in the theory of small samples. Biometrika, 22(3/4):253–283, May 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332095.

Pearson:1936:EST

[PS36] E. S. Pearson and C. Chandra Sekar. The efficiency of statistical tools and a criterion for the rejection of outlying observations. Biometrika, 28(3/4):308–320, December 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333954.

Pearson:1932:FAS

[PSD32] Karl Pearson, S. A. Stouffer, and F. N. David. Further applications in statistics of the T_{m}(x) Bessel function. Biometrika, 24 (3/4):293–350, November 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331971.

Pearson:1933:MSA

[PW33] E. S. Pearson and S. S. Wilks. Methods of statistical analysis appropriate for k samples of two variables. Biometrika, 25(3/4):353–378, December 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332289.

Pearson:1935:FIM

[PW35a] Karl Pearson and T. L. Woo. Further investigation of the morphometric characters of the individual bones of the human skull. Biometrika, 27(3/4):424–465, October 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332701.

Przyborowski:1935:SPR

[PW35b] J. Przyborowski and H. Wileński. Statistical principles of routine work in testing clover seed for dodder. Biometrika, 27(3/4):273–292, October 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332691.

Przyborowski:1936:NAT

[PW36] J. Przyborowski and H. Wilenski. Note on the application of a theorem of Frau Pollaczek-Geiringer. Biometrika, 28(1/2):187–188, June 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334130.

Roberts:1939:OTR

[RDM39] E. Roberts, W. M. Dawson, and Margaret Madden. Observed and theoretical ratios in Mendelian inheritance. Biometrika, 31(1/2):56– 66, July 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334975.

Rider:1932:DCC

[Rid32] Paul R. Rider. On the distribution of the correlation coefficient in small samples. Biometrika, 24(3/4):382–403, November 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2331973.

Rietz:1932:SNN

[Rie32] H. L. Rietz. A simple non-normal correlation surface. Biometrika, 24(1/2):288–290, May 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333802.

Rigby:1933:RRC

[Rig33] Constance M. Rigby. On a recurrence relation connected with the double Bessel functions K_{τ_{1},τ_{2}}(x) and T_{τ_{1},τ_{2}}(x). Biometrika, 25 (3/4):420–421, December 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332293.

Risdon:1939:SCO

[Ris39] D. L. Risdon. A study of the cranial and other human remains from Palestine excavated at Tell Duweir (Lachish) by the Wellcome– Marston Archaeological Research Expedition. Biometrika, 31(1/2): 99–166, July 1939. CODEN BIOKAX. ISSN 0006-3444 (print),

1464-3510 (electronic). URL http://www.jstor.org/stable/
2334978.

Romanovsky:1933:PMR

[Rom33] V. Romanovsky. On a property of the mean ranges in samples from a normal population and on some integrals of Prof. T. Hojo. Biometrika, 25(1/2):195–197, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332275.

Romanovsky:1936:NMM

[Rom36] V. Romanovsky. Note on the method of moments. Biometrika, 28(1/2):188–190, June 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334131.

S:1939:BRB

[S.39] H. L. S. Book review: Twenty-Five Years of Health Progress. Biometrika, 30(3/4):469–470, January 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332660.

Sandon:1938:ARR

[San38] Frank Sandon. Absence and recovery recorded at irregular intervals, illustrated on school data. Biometrika, 29(3/4):336–349, February 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332009.

Shenton:1979:CPR

[SBL79] L. R. Shenton, K. O. Bowman, and H. K. Lam. Comments on a paper by R. C. Geary on standardized mean deviation. Biometrika, 66(2):400–401, August 1979. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2335679. See [Gea36].

Stoessiger:1932:SCV

[SM32] Brenda N. Stoessiger and G. M. Morant. A study of the crania in the vaulted ambulatory of Saint Leonard’s Church, Hythe. Biometrika, 24(1/2):135–202, May 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333799.

StGeorgescu:1932:FCS

[St.32] N. St. Georgescu. Further contributions to the sampling problem. Biometrika, 24(1/2):65–107, May 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333797.

Steffensen:1931:SRR

[Ste31] J. F. Steffensen. Some recent researches in the theory of statistics and actuarial science. Biometrika, 22(3/4):424–425, May 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332105.

Steffensen:1934:CMD

[Ste34] J. F. Steffensen. On certain measures of dependence between statistical variables. Biometrika, 26(1/2):251–255, May 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332058.

Student:1938:CBB

[Stu38] Student. Comparison between balanced and random arrangements of field plots. Biometrika, 29(3/4):363–378, February 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332011.

Suski:1933:BBA

[Sus33] P. M. Suski. The body build of American-born Japanese children. Biometrika, 25(3/4):323–352, December 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332288.

Tang:1938:CSP

[Tan38] Y. Tang. Certain statistical problems arising in plant breeding. Biometrika, 30(1/2):29–56, June 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332222.

Thompson:1935:SVR

[TE35] Catherine M. Thompson and Ethel M. Elderton. Strength of vision in relation to eye and hair pigmentation. Biometrika, 27(1/2):266– 268, March 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332047.

Teichroew:1965:HDS

[Tei65] Daniel Teichroew. A history of distribution sampling prior to the era of the computer and its relevance to simulation. Journal of the American Statistical Association, 60(309):27–49, March 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/stable/2283135.

Terentjev:1931:BUM

[Ter31] Paul V. Terentjev. Biometrische Untersuchungen über die Morpho-Logischen Merkmale von Rana Ridibunda Pall: (Amphibia, Salientia) (German) [Biometric studies on the morphological characteristics of Rana ridibunda Pall: (Amphibia, Salientia)]. Biometrika, 23(1/2):23–51, November 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333629.

Thompson:1932:GPMb

[THM^{+}32] William R. Thompson, Raymond Hussey, Joseph T. Matteis, William C. Meredith, George C. Wilson, and F. Erwin Tracy. The geometric properties of microscopic configurations. II. Incidence and volume of Islands of Langerhans in the pancreas of a monkey. Biometrika, 24(1/2):27–38, May 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333793.

Thompson:1932:GPMa

[Tho32] William R. Thompson. The geometric properties of microscopic configurations. I. General aspects of projectometry. Biometrika, 24(1/2):21–26, May 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333792.

Thompson:1933:LOU

[Tho33] William R. Thompson. On the likelihood that one unknown probability exceeds another in view of the evidence of two samples. Biometrika, 25(3/4):285–294, December 1933. CODEN BIOKAX.

ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332286.

Tildesley:1933:ANS

[Til33] Miriam L. Tildesley. The Albanians of the North and South: (1) introductory account of measurements and photographs taken in 1929. Biometrika, 25(1/2):21–51, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332259.

Usher:1932:HEH

[Ush32] C. H. Usher. Hereditary entropion and hereditary changes in the skin of the eyelids. Biometrika, 24(1/2):1–20, May 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2333791.

Usher:1935:PHE

[Ush35] C. H. Usher. Pedigrees of hereditary epicanthus. Biometrika, 27(1/2):5–25, March 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332038.

vonBonin:1936:COC

[vB36] Gerhardt von Bonin. On the craniology of Oceania. Crania from New Britain. Biometrika, 28(1/2):123–148, June 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334127.

VonBonin:1938:IRU

[VM38] Gerhardt Von Bonin and G. M. Morant. Indian races in the United States. A survey of previously published cranial measurements. Biometrika, 30(1/2):94–129, June 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332227.

VonBonin:1931:BKO

[Von31] Gerhardt Von Bonin. Beitrag Zur Kraniologie von Ost-Asien. (German) [Contribution to the craniology of East Asia]. Biometrika, 23(1/2):52–113, November 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2333630.

Wagner:1935:EDI

[Wag35] K. Wagner. Endocranial diameters and indices. A new instrument for measuring internal diameters of the skull. Biometrika, 27(1/2):88–132, March 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332040.

Welch:1935:SPA

[Wel35] B. L. Welch. Some problems in the analysis of regression among k samples of two variables. Biometrika, 27(1/2):145–160, March 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332042.

Welch:1937:ZTR

[Wel37] B. L. Welch. On the z-test in randomized blocks and Latin squares. Biometrika, 29(1/2):21–52, June 1937. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332405.

Welch:1938:TH

[Wel38a] B. L. Welch. On tests for homogeneity. Biometrika, 30(1/2):149–158, June 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332230.

Welch:1938:SDB

[Wel38b] B. L. Welch. The significance of the difference between two means when the population variances are unequal. Biometrika, 29(3/4):350–362, February 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332010.

Welch:1939:NDF

[Wel39a] B. L. Welch. Note on discriminant functions. Biometrika, 31 (1/2):218–220, July 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2334985.

Welch:1939:DML

[Wel39b] B. L. Welch. On the distribution of maximum likelihood estimates. Biometrika, 31(1/2):187–190, July 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334981.

Wicksell:1933:CFT

[Wic33] S. D. Wicksell. On correlation functions of Type III. Biometrika, 25(1/2):121–133, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332267.

Wilks:1932:CGA

[Wil32] S. S. Wilks. Certain generalizations in the analysis of variance. Biometrika, 24(3/4):471–494, November 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331979.

Williams:1935:NSD

[Wil35] P. Williams. Note on the sampling distribution of √β_{1}, where the population is normal. Biometrika, 27(1/2):269–271, March 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332048.

Wishart:1930:DCH

[Wis30] John Wishart. The derivation of certain high order sampling product moments from a normal population. Biometrika, 22(1/2):224–238, July 1930. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2332072.

Wishart:1932:NDC

[Wis32] John Wishart. A note on the distribution of the correlation ratio. Biometrika, 24(3/4):441–456, November 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331977.

Wishart:1933:CSI

[Wis33] John Wishart. A comparison of the semi-invariants of the distributions of moment and semi-invariant estimates in samples from an infinite population. Biometrika, 25(1/2):52–60, May 1933. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332260.

Wisniewski:1935:VCP

[Wiś35] Jan Wiśniewski. On the validity of a certain Pearson’s formula. Biometrika, 27(3/4):356–363, October 1935. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2332696. See note [Pea35b].

Wisniewski:1936:VCP

[Wis36] Jan Wisniewski. On the validity of a certain Pearson’s formula — a rejoinder. Biometrika, 28(1/2):190–192, June 1936. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334132.

Wishart:1938:GRD

[Wis38] John Wishart. Growth-rate determinations in nutrition studies with the bacon pig, and their analysis. Biometrika, 30(1/2):16–28, June 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2332221.

Wishart:1931:MSM

[WKE31] J. Wishart, T. Kondo, and E. M. Elderton. The mean and second moment coefficient of the multiple correlation coefficient, in samples from a normal population. Biometrika, 22(3/4):353–376, May 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332101.

Woo:1932:PCA

[WM32] T. L. Woo and G. M. Morant. A preliminary classification of Asiatic races based on cranial measurements. Biometrika, 24(1/2):108–134, May 1932. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333798.

Woo:1934:BSF

[WM34] T. L. Woo and G. M. Morant. A biometric study of the “flatness” of the facial skeleton in man. Biometrika, 26(1/2):196–250, May 1934. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332057.

Woo:1930:SSO

[Woo30] T. L. Woo. A study of seventy-one Ninth Dynasty Egyptian skulls from Sedment. Biometrika, 22(1/2):65–93, July 1930. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332069.

Woo:1931:AHS

[Woo31] T. L. Woo. On the asymmetry of the human skull. Biometrika, 22(3/4):324–352, May 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332100.

Woo:1937:BSH

[Woo37] T. L. Woo. A biometric study of the human malar bone. Biometrika, 29(1/2):113–123, June 1937. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332409.

Wilks:1937:SDC

[WT37] S. S. Wilks and Catherine M. Thompson. The sampling distribution of the criterion λ_{H_{1}} when the hypothesis tested is not true. Biometrika, 29(1/2):124–132, June 1937. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332410.

Wunderly:1939:COS

[Wun39] J. Wunderly. The cranial and other skeletal remains of tasmanians in collections in the Commonwealth of Australia. Biometrika, 30(3/4):305–337, January 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2332652.

Yates:1939:CAS

[Yat39] F. Yates. The comparative advantages of systematic and randomized arrangements in the design of agricultural and biological experiments. Biometrika, 30(3/4):440–466, January 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332658.

Young:1931:WSS

[You31] Matthew Young. The West Scottish skull and its affinities. Biometrika, 23(1/2):10–22, November 1931. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2333628.

Young:1938:LS

[You38] Matthew Young. The London skull. Biometrika, 29(3/4):277–321, February 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332007.

Yule:1938:NKP

[Yul38a] G. U. Yule. Notes of Karl Pearson’s lectures on the theory of statistics, 1884–96. Biometrika, 30(1/2):198–203, June 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332238.

Yule:1938:SPN

[Yul38b] G. Udny Yule. On some properties of normal distributions, univariate and bivariate, based on sums of squares of frequencies. Biometrika, 30(1/2):1–10, June 1938. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332219.

Yule:1939:SLS

[Yul39] G. Udny Yule. On sentence-length as a statistical characteristic of style in prose: With application to two cases of disputed authorship. Biometrika, 30(3/4):363–390, January 1939. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2332655.